# 1st Chenchen Wang College of Information Science

### v1

@@ Abstract
This paper develops a reinforcement learning-based controller for the stabilization of partial differential equation (PDE) systems. Within the soft actor-critic (SAC) framework, we embed a DeepONet, a well-known neural operator (NO), which is pretrained using the backstepping controller. The pretrained DeepONet captures the essential features of the backstepping controller and serves as a feature extractor, replacing the convolutional neural networks (CNNs) layers in the original actor and critic networks, and directly connects to the fully connected layers of the SAC architecture. We apply this novel backstepping and reinforcement learning integrated method to stabilize an unstable first-order hyperbolic PDE and an unstable reaction-diffusion PDE. Simulation results demonstrate that the proposed method outperforms the standard SAC, SAC with an untrained DeepONet, and the backstepping controller on both systems.
Index Terms:
Reinforcement Learning, soft actor-critic, DeepONet, first-order hyperbolic PDEs, diffusion-reaction PDEs.
@@ I Introduction
For finite-dimensional systems, there has been a growing body of research [2, 3, 4]. For instance, a control-informed reinforcement learning framework that integrates PID control components into the architecture of deep reinforcement learning (RL) policies is proposed in [5], In this [6], it embeds barrier function theory into neural ordinary differential equations (ODEs), ensuring that the system state remains within a safe region by optimizing the parameters of a learning-based barrier function.
@@
In our method, the DeepONet pretrained on the backstepping controller serves as a feature extractor, replacing the CNN layers in the standard SAC architecture. The DeepONet connects directly to the fully connected layers of the actor and critic networks. During training, the parameters of the DeepONet feature extractor are updated alongside the SAC actor and critic networks, enabling fine-tuning in the RL process.
@@
The structure of this paper is as follows: Section II provides problem formulation; Section III presents the NO-based reinforcement learning framework; Section IV validates the effectiveness of the method through simulations; and finally, Section V concludes the paper.
@@ II-A General PDE control problem
with the initial condition $u(x,0)=u_{0}(x)$, where $F$ is the PDE that models the system dynamics, $U(t)$ is the control function, and $p_{i}(x)$, for $i=1,\dots,m$ are the coefficient functions with $m$ coefficients. The PDE control problem is to design a controller which stabilize the system or a cost function regulates $u(x,t)$ to a desired trajectory.
@@ II-B Backstepping control operator
Employing the PDE backstepping method, one can design a feedback controller, usually the boundary controller, which is generalized by
[…]
where for given $p_{i}(x)\in C[0,1]$ and $u\in C[0,1]$,
[…]
with $k(x,y)$ denoting the backstepping kernel function dependent on the system coefficients.
@@ II-C DeepONet for learning backstepping controller
DeepONet is a deep learning architecture specifically designed for learning “function-to-function” mappings (operators). It consists of two subnetworks: a branch net and a trunk net. The output function is expressed as a weighted sum of basis functions, where the branch net learns the mapping from the input function to the coefficients (which serve as the weights of the basis functions), and the trunk net learns the basis functions themselves.
For the PDE control problem, the inputs to the DeepONet consist of the coefficients $p_{i}(x)$ and state $u(x,\cdot)$, while the output of the DeepONet is the control signal $U(\cdot)$. All training data are generated by solving the PDE (1) with the backstepping controller (2), using different initial conditions and coefficient functions randomly sampled from Chebyshev polynomials with random parameters [11].
The trained DeepONet is adapted for integration into the SAC framework by modifying its output from a control value to a feature vector, as shown in Fig. 1. This feature vector serves as input to the FNN layers of both the actor and critic networks, introducing the backstepping control knowledge into the RL learning.
@@ III NO-based reinforcement learning framework
In this section, we present the RL framework that integrates a pre-trained DeepONet with the SAC algorithm. We use a SAC framework with an actor network $\pi_{\phi}(a_{t}|s_{t})$, two critic networks $Q_{\theta_{i}}(s_{t},a_{t})$, and their corresponding target networks $Q_{\bar{\theta}_{i}}(s_{t},a_{t})$ (the dual-critic design mitigates bootstrapping issues). The DeepONet introduced in Section II-C, is pre-trained using the backstepping controller, which serves as a feature extractor and replaces the CNN layers in the standard SAC architecture. The DeepONet connects directly to the actor/critic fully connected layers. During training, the actor and critic networks, along with their embedded DeepONet feature extractors, are jointly trained via backpropagation. The architecture of the RL is shown in Fig. 2.
@@ III-A Actor network
By introducing the policy entropy $H_{\pi}$, it measures the randomness of the action distribution given the state $s_{t}$. The policy entropy is defined as: $H(\pi(\cdot|s_{t}))=-\mathbb{E}_{a_{t}\sim\pi_{\phi}}[\log\pi_{\phi}(a_{t}|s_{t})]$. The policy network $\pi_{\phi}(a_{t}|s_{t})$ is trained by minimizing the following objective function:
$J_{\pi}({\phi})=\mathbb{E}_{s_{t}\sim\mathcal{G}}\mathbb{E}_{a_{t}\sim\pi_{\phi}}\left[-\alpha H(\pi_{\phi}(\cdot|s_{t}))-\min_{i=1,2}Q_{\theta_{i}}(s_{t},a_{t})\right],$ (4)
where $\mathcal{G}$ denotes the state distribution. Accordingly, the parameter $\phi$ of the policy network is updated by
$\phi\leftarrow\phi-\eta\nabla_{\phi}J_{\pi}(\phi),$ (5)
where $\eta$ denotes the learning rate.
@@ III-B Critic networks
We use a double Q-network structure (two independent FNNs), the action-value function is defined as $Q_{\pi}(s_{t},a_{t})$, which represents the expected cumulative return that the agent obtains by following the policy $\pi_{\phi}$ after executing the action $a_{t}$ in the state $s_{t}$. Using two target networks $Q_{\bar{\theta}_{1,2}}$ to estimate the target action-value:
@@
$\alpha$ is the temperature parameter, which is used to balance the trade-off between reward maximization and policy entropy, and the action $\hat{a}_{t+1}\sim\pi(\cdot|s_{t+1})$ is sampled from the policy network. The weights of the target network $Q_{\bar{\theta}_{i}}$ are updated by
@@
where $i\in\{1,2\}$ and $\mathcal{W}$ is the distribution of the states and actions. The parameter $\theta_{i}$ is updated via stochastic gradient descent:
$\theta_{i}\leftarrow\theta_{i}-\lambda\hat{\nabla}_{\theta_{i}}{J}_{Q}(\theta_{i}),$ (9)
where $\lambda$ denotes the learning rate.
@@ III-C Algorithm (Markov Decision Process)
1) State space $\mathcal{S}$: we denote $s_{t}=\{u(x,t)\}\in\mathcal{S}$ as the augmented state at time $t$.
2) Action space $\mathcal{A}\subseteq\mathbb{R}$: it is defined as $\mathcal{A}=[-\bar{U},\bar{U}]$, where $\bar{U}=\sup_{t\in\mathbb{R}^{+}}|U(t)|$, $U(t)=a_{t}\in\mathcal{A}$.
3) Policy function $\pi(a_{t}|s_{t})$: the policy is a probability density function $\pi(a_{t}|s_{t})$ that maps the current state $s_{t}\in\mathcal{S}$ to a probability distribution over actions $a_{t}\in\mathcal{A}$.
4) State transition $p(s_{t+1}|s_{t},a_{t})$: this $p(s_{t+1}|s_{t},a_{t})$ is the probability distribution of $s_{t+1}$, which depends on the current state $s_{t}$ and the action $a_{t}$ chosen by the agent. Given the state $s_{t}$ at time $t$, when the agent executes action $a_{t}$, the environment transitions to the next state $s_{t+1}$ following this probability distribution.
@@
$q(s_{T},a_{0},\ldots,a_{T})$ =$=\begin{cases}0&\|s_{T}\|_{L_{2}}>\zeta,\\ \begin{aligned} \sigma-\dfrac{1}{\eta}\cdot\sum_{\tau=0}^{T}|a_{\tau}|_{L_{1}}\\ -\|s_{t}\|_{L_{2}}\end{aligned}&\|s_{T}\|_{L_{2}}\leq\zeta,\end{cases}$ (11)
here, $\sigma$, $\eta$, and $\zeta$ are hyperparameters.
The reward function $r_{t}$ is applied at each step to drive state convergence, while an additional reward $q$ is given at the end of each episode if the $L_{2}$ norm of the final state is less than or equal to a specified threshold $\zeta$.
Algorithm 1 SAC Procedure
1:Initialization: $\phi,\theta_{i},\bar{\theta}_{i}$ for $i=1,2$, $\mathcal{W}$ and $s_{0}=\{u_{0}(x)\}$
2:for $l\leftarrow 0$ to total steps do
3:     if $\lVert s_{t}\rVert_{L_{2}}>\mathrm{limit}\lor t\geq T$ then
4:         Reset initial state $s_{0}$
@@
6:     Observe current state $s_{t}=\{u_{t}\}$
7:     Sample action $a_{t}\sim\pi_{\phi}(a_{t}\mid s_{t})$
8:     Execute $a_{t}$, observe $s_{t+1}$ and $r_{t}$
9:     $\mathcal{W}\leftarrow\mathcal{W}\cup\{(s_{t},a_{t},r_{t},s_{t+1})\}$
10:     for each gradient step do
11:         Sample a mini-batch from $\mathcal{W}$
12:         Compute the target value $y_{t}$ using (6)
13:         Update parameters of the $Q_{\theta_{1,2}}$ by (8)
14:         Update parameters of the $Q_{\bar{\theta}_{1,2}}$ by (5)
15:         Update parameters of the $\pi_{\phi}$ by (4)
@@ IV Simulation
In this section, we apply the proposed RL controller to both a hyperbolic PDE and a parabolic PDE. Details of the corresponding backstepping designs can be found in [19]. The standard SAC code is obtained from the PDE Control Gym platform introduced in [20]. We compare our method (NOSAC_training) with the backstepping controller (Backstepping), the standard SAC (SAC), and SAC embedded with DeepONet without pretraining the DeepONet using the backstepping controller (NOSAC).
Note that all the experiments are carried out on a workstation equipped with an Intel Core i9-13900KF CPU and an NVIDIA GeForce RTX 4090 GPU. For all the hyperparameters required by the SAC algorithm and the hyperparameters of the reward function, please refer to Appendix A.2 of [20].
@@ IV-A Simulation of a 1D Hyperbolic PDE
First, we train a DeepONet to approximate the backstepping controller. To generate the training data. We employ the finite difference method with a spatial step size of $dx=0.01$ over the time interval $[0,5]$ with a temporal step size of $\Delta t=0.001$. The coefficient $\beta(x)$ is sampled from the Chebyshev polynomial $\beta(x)=5\cos\big{(}\gamma\cos^{-1}(x)\big{)}$, where $\gamma\sim U[5.5,7]$. The initial condition $u_{0}(x)$ is assumed to be a constant randomly chosen from the uniform distribution $U[1,10]$.
We generate a dataset comprising $6\times 10^{5}$ samples, which were produced by combining 100 different $\beta(x)$ and 60 different $u_{0}(x)$. During the data collection process, data was collected every 50 time steps. After completing the entire data collection, we randomly shuffled the dataset and split it into a training set 90% and a testing set 10%.
@@
The DeepONet approximation results are shown in Fig. 3, illustrating the boundary control $U(t)$ and the $L_{2}$ norm of the state under the backstepping controller and the NO-based controller for $\gamma=5.5$. It can be seen that the state converges to zero under both controllers.
@@
The same $\gamma$ value is used, and the initial state $u_{0}(x)$ is randomly selected during RL. Training consists of 100,000 steps, with 100 interaction steps per episode. Fig. 4(a) shows the reward curves for SAC, NOSAC, and NOSAC_training on 1D hyperbolic PDE, where NOSAC_training achieves faster reward growth and finds effective policies more quickly than the other two.
We apply the NOSAC_training controller to the 1D hyperbolic PDE and compare its performance with the backstepping, SAC, and NOSAC, as shown in Fig. 5. The backstepping controller yields smoother closed-loop state evolution, while the proposed NOSAC_training controller exhibits less overshoot. To better illustrate performance, Fig. 6 plots the control effort and the $L_{2}$ norm of the state, showing that NOSAC_training outperforms the backstepping controller, SAC, and NOSAC in terms of overshoot and convergence rate, while the backstepping controller achieves lower steady-state error.
To evaluate robustness under model mismatch, we simulate a system with $\gamma=5.7$ using the three LR_based controllers trained at $\gamma=5.5$ and the backstepping controller designed for $\gamma=5.5$. The results in Fig. 6 demonstrate that NOSAC_training achieves superior robustness, outperforming the other controllers in overshoot, convergence rate, and steady-state error.

### v2

# 1st Chenchen Wang    2nd Jie Qi    3rd Jiaqi Hu

@@ Abstract
This paper develops a reinforcement learning-based controller for the stabilization of partial differential equation (PDE) systems. Within the Soft Actor-Critic (SAC) framework, we embed a DeepONet, a well-known neural operator (NO), which is pretrained using the backstepping controller. The pretrained DeepONet captures the essential features of the backstepping controller and is used to extract features, replacing the convolutional neural networks (CNNs) in the original actor-critic networks, and directly connects to the fully connected layers of the SAC architecture. We apply this novel backstepping and reinforcement learning integrated method to stabilize an unstable 1D hyperbolic PDE and an unstable reaction-diffusion PDE. The proposed method is shown to outperform standard SAC in simulations, SAC with an untrained DeepONet, and the backstepping controller on both systems.
@@ I Introduction
For finite-dimensional systems, there has been a growing body of research [2, 3, 4]. For instance, a control-informed reinforcement learning (RL) framework that integrates PID control components into the architecture of deep RL policies is proposed in [5], In this [6], it embeds barrier function theory into neural ordinary differential equations (ODEs), ensuring that the system state remains within a safe region by optimizing the parameters of a learning-based barrier function.
@@
In our method, the DeepONet pretrained on the backstepping controller is used for feature extraction, as a replacement for the CNNs in standard SAC. The DeepONet connects directly to the fully connected actor-critic networks layers. Throughout the training phase, parameters of the DeepONet feature extraction module are optimized in conjunction with those of the SAC actor-critic networks, thereby enabling fine-tuning during reinforcement learning.
@@
The structure of this paper is as follows: Section II provides problem formulation; Section III presents the NO-based RL framework; Section IV validates the effectiveness of the method through simulations; and finally, Section V concludes the paper.
@@ II-A General PDE control problem
with the initial condition $u(x,0)=u_{0}(x)$, where $F$ is the PDE that models the system dynamics, $U(t)$ is the control function, and $p_{i}(x)$, for $i=1,\dots,m$ are the coefficient functions. The PDE control problem is to design a controller which stabilizes the system or regulates states to a desired trajectory.
@@ II-B Backstepping control operator
Employing the PDE backstepping method [19], a feedback controller—typically implemented as a boundary controller—can be designed in the following form:
$U(t)$ $=\int_{0}^{1}k(1,y)u(y,t)\,dy,$ (2)
where $k(1,y)$ is the control gain derived from the backstepping kernel function. This kernel depends on the system coefficient functions $p_{i}(x)$, $i=1,\ldots,m$. To formalize the controller structure, we define the control operator $\mathcal{U}:(C[0,1])^{m}\times L^{2}[0,1]\mapsto\mathbb{R}$, which maps the coefficient functions $(p_{1},\ldots,p_{m})\in(C[0,1])^{m}$ and the system state $u\in L^{2}[0,1]$ to the control input $U$. Specifically,
$U:=\mathcal{U}(p_{1},\dots,p_{m},u),$ (3)
where $U$ is defined as in (2).
@@ II-C DeepONet for learning backstepping controller
DeepONet is a deep learning architecture designed to learn function-to-function mappings, consisting of branch and trunk networks. The output function is a weighted sum of basis functions, with the branch network learning the weights and the trunk network learning the basis functions.
For the PDE controller approximated by a DeepONet, the DeepONet receives coefficient functions $p_{i}(x)$ and state $u(x,\cdot)$ as inputs, generating a scalar control signal $U(\cdot)$ as output. In the mapping, the input and output are time-independent, the control input is generated based on the current system state. All training data are generated by solving the PDE (1) with the backstepping controller (2), using different initial conditions and coefficient functions randomly sampled from Chebyshev polynomials with random parameters [11].
To train the DeepONet, we employ the following state-based loss function:
$\varepsilon=|U_{NO}-U|,$ (4)
where $U_{NO}$ denotes the control generated by DeepONet, while $U$ denotes the backstepping control.
To integrate the DeepONet into the reinforcement learning architecture, we modify its output from a scalar control signal to a vector representation, enabling feature extraction for both actor-critic networks in SAC, as depicted in Fig. 1. The dimensionality of the DeepONet output is determined by its branch and trunk networks, along with subsequent fully connected layers. The pre-trained DeepONet encapsulates the knowledge of the backstepping controller, enabling seamless integration of prior control strategies into the RL decision-making process. This incorporation not only improves the training efficiency of the reinforcement learning algorithm but also enhances the performance of the RL-based controller.
@@ III NO-based reinforcement learning framework
In this section, we elaborate on the reinforcement learning (RL) framework that fuses a pre-trained DeepONet with the Soft Actor-Critic (SAC) algorithm. This framework is composed of an actor network $\pi_{\phi}(a_{t}|s_{t})$, a pair of critic networks denoted as $Q_{\theta_{i}}(s_{t},a_{t})$, as well as their respective target networks $Q_{\bar{\theta}_{i}}(s_{t},a_{t})$. Serving as a feature extraction component, the DeepONet replaces the CNNs adopted in the canonical SAC structure, and establishes a direct connection with the fully connected layers of the actor and critic networks. In the training phase, the actor-critic networks and the embedded DeepONet feature extractor are optimized in a joint manner through the backpropagation algorithm. The schematic diagram of this RL framework is presented in Fig. 2.
@@ III-A Actor network
By introducing the policy entropy $H_{\pi}$, it quantifies the uncertainty of the action distribution conditioned on state $s_{t}$. The policy entropy is defined as: $H(\pi(\cdot|s_{t}))=-\mathbb{E}_{a_{t}\sim\pi_{\phi}}[\log\pi_{\phi}(a_{t}|s_{t})]$. The policy network $\pi_{\phi}(a_{t}|s_{t})$ is trained by optimizing the following subsequent objective function:
$J_{\pi}({\phi})=\mathbb{E}_{s_{t}\sim\mathcal{D}}\mathbb{E}_{a_{t}\sim\pi_{\phi}}\left[-\alpha H(\pi_{\phi}(\cdot|s_{t}))-\min_{i=1,2}Q_{\theta_{i}}(s_{t},a_{t})\right],$ (5)
with $\mathcal{D}$ characterizes the distribution of system states. Consequently, the update of the policy network parameter $\phi$ is realized by
$\phi\leftarrow\phi-\eta\nabla_{\phi}J_{\pi}(\phi).$ (6)
@@ III-B Critic networks
We use a double Q-network structure (two independent FNNs), the action-value function is defined as $Q_{\pi}(s_{t},a_{t})$, which represents the expected cumulative return that the agent obtains by following the policy $\pi_{\phi}$ after executing the action $a_{t}$ in the state $s_{t}$. Using two target networks $Q_{\bar{\theta}_{1,2}}$ to determine the target Q-function:
@@
$\alpha$ is the temperature parameter for balancing reward maximization and policy entropy, and the action $\hat{a}_{t+1}\sim\pi(\cdot|s_{t+1})$ is drawn from the actor network. The weights of the target network $Q_{\bar{\theta}_{i}}$ are updated by
@@
where $i\in\{1,2\}$ and $\mathcal{W}$ is the distribution of the states and actions. The parameter $\theta_{i}$ is updated by stochastic gradient descent:
$\theta_{i}\leftarrow\theta_{i}-\lambda\hat{\nabla}_{\theta_{i}}{J}_{Q}(\theta_{i}).$ (10)
@@ III-C Algorithm (Markov Decision Process)
1) State space $\mathcal{S}$: we denote $s_{t}=\{u(x,t)\}\in\mathcal{S}$, which constitutes the extended state at time $t$.
2) Action space $\mathcal{A}\subseteq\mathbb{R}$: it is $\mathcal{A}=[-\bar{U},\bar{U}]$, with $\bar{U}=\sup_{t\in\mathbb{R}^{+}}|U(t)|$, $U(t)=a_{t}\in\mathcal{A}$.
3) Policy function $\pi(a_{t}|s_{t})$: this policy is formulated as a probability density function $\pi(a_{t}|s_{t})$, which establishes a mapping relationship from the current state space $\mathcal{S}$ to the probability distribution over the action space $\mathcal{A}$.
4) State transition $p(s_{t+1}|s_{t},a_{t})$: this $p(s_{t+1}|s_{t},a_{t})$ is the probability distribution of $s_{t+1}$, which depends on the current state $s_{t}$ and the action $a_{t}$ chosen by the agent.
@@
$q(s_{T},a_{0},\ldots,a_{T})$ $=\begin{cases}0&\|s_{T}\|_{L_{2}}>\zeta,\\ \begin{aligned} \sigma-\dfrac{1}{\eta}\cdot\sum_{\tau=0}^{T}|a_{\tau}|_{L_{1}}\\ -\|s_{t}\|_{L_{2}}\end{aligned}&\|s_{T}\|_{L_{2}}\leq\zeta,\end{cases}$ (12)
here, $\sigma$, $\eta$, and $\zeta$ are hyperparameters. The reward function $r_{t}$ drives state convergence at each step, with an additional reward $q$ given at the end of each episode if the $L_{2}$ norm of the final state is less than or equal to the threshold $\zeta$.
Algorithm 1 DeepONet-SAC Training Process
1:Initialize network parameters: $\phi,\theta_{i},\bar{\theta}_{i}$ ($i=1,2$), replay memory $\mathcal{W}$, and starting state $s_{0}=\{u_{0}(x)\}$
2:for iteration $l$ from 0 to max steps do
3:  if $\|s_{t}\|_{L_{2}}>\text{threshold}$ or $t\geq T$ then
4:   Reinitialize state $s_{t}\leftarrow s_{0}$
@@
6:  Set the instantaneous state $s_{t}=\{u_{t}\}$
7:  Draw action $a_{t}\sim\pi_{\phi}(\cdot\mid s_{t})$
8:  Apply $a_{t}$, observe next state $s_{t+1}$ and reward $r_{t}$
9:  Store transition $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{W}$
10:  for each gradient update iteration do
11:   Randomly sample mini-batch from $\mathcal{W}$
12:   Calculate target value $y_{t}$ via equation (6)
13:   Adjust critic parameters $Q_{\theta_{1,2}}$ using (8)
14:   Update target network parameters $Q_{\bar{\theta}_{1,2}}$ by (5)
15:   Modify actor parameters $\phi$ according to (4)
@@ IV Simulation
In this section, the proposed RL controller is applied to both hyperbolic and parabolic PDEs, and the details of the backstepping design can be found in [19]. We adopt the standard SAC implementation from the PDE Control Gym platform introduced in [20]. Our proposed method—SAC embedded with DeepONet pre-trained using the backstepping controller (referred to as NOSAC_training)—is compared against four baselines: the backstepping controller (Backstepping), standard SAC (SAC), and SAC embedded with DeepONet without pretraining (NOSAC).
Experiments conducted on an Intel i9-13900KF/RTX 4090 workstation; SAC and reward hyperparameters in [20] Appendix A.2.
@@ IV-A Simulation of a 1D Hyperbolic PDE
First, A DeepONet is trained to emulate the backstepping controller, with corresponding training data generated accordingly. We employ the finite difference method with $dx=0.01$ over the time interval $[0,5]$ with a temporal step size of $\Delta t=0.001$. The coefficient $\beta(x)$ is sampled from the Chebyshev polynomial $\beta(x)=5\cos\big(\gamma\cos^{-1}(x)\big)$, where $\gamma\sim U[5.5,7]$. The initial condition $u_{0}(x)$ is assumed to be a constant randomly chosen from the uniform distribution $U[1,10]$.
We generate a dataset of $6\times 10^{5}$ samples, which comprises 100 different $\beta(x)$ and 60 different $u_{0}(x)$. We collect data every 50 time steps, and then randomly shuffle and split the dataset into a training set 90% and a testing set 10%. The training process requires approximately 175 minutes.
The DeepONet approximation results are shown in Fig. 3, illustrating the boundary control $U(t)$ and the $L_{2}$ norm of the state under the backstepping controller and the NO-based controller for $\gamma=5.5$. It can be seen that $U_{NO}$ remains almost indistinguishable from $U$, and the state converges to zero under both controllers ($\|u_{NO}-u\|_{L_{2}}$ is less than $10^{-4}$ after about $3$ seconds).
@@
In the RL training, we set $\gamma=5.5$ and sample $u_{0}(x)$ sampled from a uniform distribution $U[1,10]$ at each iteration. The training consists of 100,000 steps with 100 interaction steps per episode. The training time is about 18, 20, and 11 minutes for NOSAC_training, NOSAC, and SAC taking, respectively. Fig. 4(a) shows NOSAC_training achieves faster reward growth and policy convergence compared to the others, thanks to the pre-trained DeepONet. Pre-trained on backstepping controller, the DeepONet captures stable control behaviors, enabling it to encode this knowledge into feature vectors with higher fidelity. This serves as a warm start, for the SAC actor—reducing exploration costs and increasing initial rewards—and provides the critic with more informative state features.
We apply the NOSAC_training controller to the 1D hyperbolic PDE and compare its performance with the backstepping, SAC, and NOSAC, as shown in Fig. 5. The backstepping controller yields smoother closed-loop state evolution, while the proposed NOSAC_training controller exhibits less overshoot. To better illustrate performance, Fig. 6 plots the control effort and the $L_{2}$ norm of the state, showing that NOSAC_training outperforms the other controllers in terms of overshoot and convergence rate, whereas the backstepping controller demonstrates superior steady-state precision.
To evaluate robustness under model mismatch, we simulate a system with $\gamma=5.7$ using the three LR_based controllers trained at $\gamma=5.5$ and the backstepping controller designed for $\gamma=5.5$. The results in Fig. 7 demonstrate that NOSAC_training achieves superior robustness, outperforming the SAC and NOSAC controllers in overshoot, convergence speed, and steady-state error.